OpenStudy (anonymous):

f(X)=-4tan4pi(X+4)-2, what is the vertical displacement and phase shift?

OpenStudy (anonymous):

Given some function \(f(x)\), a phase shift of \(c\) is given by \(f(x-c)\) and a vertical displacement of \(d\) is given by \(f(x)+d\). In this case, \(f(x)=-4\tan 4\pi x\). Note that \(f(x-c)=-4\tan4\pi(x-c)\). This means that you'll have to distribute the \(4\pi\): \[f(x-c)=-4\tan(4\pi x-4\pi c)\]